# Impact of phylogeny on the inference of functional sectors from protein sequence data

### v1

@@ Abstract
Statistical analysis of multiple sequence alignments of homologous proteins has revealed groups of coevolving amino acids called sectors. These groups of amino-acid sites feature collective correlations in their amino-acid usage, and they are associated to functional properties. Modeling showed that natural selection on an additive functional trait of a protein is generically expected to give rise to a functional sector. These modeling results motivated a principled method, called ICOD, which is designed to identify functional sectors, as well as mutational effects, from sequence data. However, a challenge for all methods aiming to identify sectors from multiple sequence alignments is that correlations in amino-acid usage can also arise from the mere fact that homologous sequences share common ancestry, i.e. from phylogeny. Here, we generate controlled synthetic data from a minimal model comprising both phylogeny and functional sectors. We use this data to dissect the impact of phylogeny on sector identification and on mutational effect inference by different methods. We find that ICOD is most robust to phylogeny, but that conservation is also quite robust. Next, we consider natural multiple sequence alignments of protein families for which deep mutational scan experimental data is available. We show that in this natural data, conservation and ICOD best identify sites with strong functional roles, in agreement with our results on synthetic data. Importantly, these two methods have different premises, since they respectively focus on conservation and on correlations. Thus, their joint use can reveal complementary information.
@@ Introduction
Statistical inference from genome and protein sequence data is currently making great progress. This is due both to the important growth of available genome sequences, and to the development of inference methods ranging from interpretable models inspired by statistical physics and information theory to large deep learning models. One important type of data sets consists in multiple sequence alignments (MSAs) of homologous proteins. These proteins, which are said to be the members of a protein family, share a common ancestry and common functional properties. Because of this, MSA columns (i.e. residue sites) feature correlations in amino-acid usage. Pairwise correlations due to contacts between amino acids in the three-dimensional structure of proteins have been thoroughly studied [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23]. Statistical analyses of MSAs have also revealed the existence of groups of collectively correlated amino acids, termed sectors, which have are associated to a functional role and are often spatially close in the protein structure [24, 25, 26, 27, 28, 29]. Preserving the pairwise correlations of natural sequences in a sector-based analysis enabled the successful design of new functional sequences in a pioneering work [26]. Sectors can be identified using Statistical Coupling Analysis (SCA), which focuses on the large-eigenvalue modes of a covariance matrix weighted by conservation [27, 30]. It was recently shown in a minimal model that sectors of collectively correlated amino acids emerge from nonlinear selection on an additive trait (i.e. functional property) of the protein [31]. This case should be quite generic, since empirical data can often be modeled as nonlinear functions of a linear trait [32], and since many protein traits involve additive contributions from amino acids [33, 34, 32]. These modeling results motivated a principled method, called ICOD (for Inverse Covariance Off-Diagonal), which is designed to identify functional sectors, and more generally mutational effects, from sequence data [31]. Here, the focus is on the large-eigenvalue modes of the ICOD matrix, which essentially correspond to the small-eigenvalue modes of the covariance matrix, making ICOD a priori quite distinct from SCA. Another important difference is that ICOD was designed to focus on correlations and eliminate conservation signal as well as possible. Accordingly, a strength of ICOD is its robustness to conservation [31].
@@ Results
To incorporate phylogeny, we take an equilibrium sequence and we evolve this ancestral sequence along a binary branching tree, see Fig 1B. Along the tree, random mutations changing the state of an amino acids are proposed, and they are accepted or rejected according to the same Metropolis criterion as the one used to generate equilibrium sequences. Thus, selection on the trait is maintained through the phylogeny. A fixed number $\mu$ of accepted mutations are performed on each branch. Generated sequences at the leaves of the tree therefore contain correlations from phylogeny and from selection. Their importance are controlled respectively by the parameters $\mu$ and $\kappa$: a smaller $\mu$ means that sequences are more closely related, yielding more phylogenetic correlations, while a larger $\kappa$ means stronger selection, yielding more correlations arising from the sector. Note that this data generation process with phylogeny is close to the one we used previously [47, 40], but that the Hamiltonian we use here (Eq 1) is specific to the sector model.
Figure 1: Data generation process. A: Generation of sequences with selection only. Given a vector of mutational effects $\vec{D}$ on the trait of interest, we sample independent equilibrium sequences under the Hamiltonian $H(\vec{\sigma})$ in Eq 1. For this, we start from random sequences and we use a Metropolis Monte Carlo algorithm where proposed mutations (changes of state at a randomly chosen site) are accepted with probability $p$, according to the Metropolis criterion associated to $H(\vec{\sigma})$. The obtained sequences feature pairwise correlations and conservation arising from selection on the trait (via the Hamiltonian $H(\vec{\sigma})$ in Eq 1). B: To incorporate phylogenetic correlations, we start from one equilibrium sequence, which becomes the ancestor. We evolve it on a binary branching tree over a fixed number of “generations” (branching events), here 2 generations. A fixed number of mutations $\mu$ (red, here $\mu=2$) are accepted with probability $p$ on each branch of the tree. The earliest generation at which a site mutates with respect to its ancestral state is denoted by $G$, see examples highlighted in green.
@@
Before studying the effect of phylogeny, let us analyze the ICOD matrix and its the spectrum without phylogeny, which will serve as baseline. To first order in $\kappa$ (small selection regime), the off-diagonal elements of the ICOD matrix can be approximated as $\tilde{C}_{ij}^{-1}\approx\kappa D_{i}D_{j}$ for all $i,j$ between 1 and $L$ with $i\neq j$ [31]. Here, we show using two different methods that this formula still holds beyond first order in $\kappa$: one method extents this formula to second order, and the second one to fourth order (see Supplementary material). Within this robust approximation, if diagonal elements were equal to $\kappa D_{i}^{2}$ instead of 0, the spectrum of the ICOD matrix would comprise a single nonzero eigenvalue $\kappa\sum_{i=1}^{L}D_{i}^{2}$ associated to the mutational effect vector $\vec{D}$. Hence, the eigenvector associated to the largest eigenvalue of the ICOD matrix allows to recover $\vec{D}$, as observed in [31].
@@
Note that in the example shown in Fig 2, important mutational effects corresponding to the sector are all positive. We checked that our results are robust to the case where the sector comprises both sites with large positive and and large negative mutational effects: the spectrum is very similar to that in Fig 2, and the top eigenvector of the ICOD matrix successfully recovers $\vec{D}$.
Figure 2: Spectrum of the ICOD matrix and of its block diagonal approximation. Left (resp. middle) panel: spectrum of the ICOD matrix $\tilde{C}^{-1}$ computed on 2000 (resp. 14,000) sequences generated independently at equilibrium, and of its block diagonal approximation. The spectrum of the inverse covariance matrix $C^{-1}$ is also shown as a reference. Right panel: spectrum of the analytical approximation of the ICOD matrix, and of its block diagonal approximation. Sequences of length $L=200$ were sampled independently at equilibrium using the Hamiltonian in Eq 1 with $\kappa=10/\left(\sum_{i}{D_{i}^{2}}\right)$ and $\tau^{*}=90$. The vector of mutational effect $\vec{D}$ comprises sector sites (the 20 first sites) with components sampled from a Gaussian distribution with mean 5 and variance 0.25, and non-sector sites (the remaining 180 sites) with components sampled from a Gaussian distribution with mean 0.5 and variance 0.25. The analytical approximation $\tilde{C}^{-1}_{ij}\approx(1-\delta_{ij})\kappa D_{i}D_{j}$ (see Supplementary material) was computed from the values of $\kappa$ and $\vec{D}$ used for data generation.
@@
Figure 3: Impact of phylogeny and selection on ICOD, covariance and SCA spectra. Eigenvalues of the ICOD, covariance and SCA matrices, sorted from largest to smallest, are shown for sequences generated with only phylogeny (light shades) and both phylogeny and selection (dark shades). We consider different levels of phylogeny by considering different values of $\mu$ (shown as different colors). ‘No phylogeny’ corresponds to sequences generated independently at equilibrium, and thus containing only correlations due to selection. This data set comprises $M=2048$ sequences generated exactly as in Fig. 2. Data sets without selection are generated by evolving random sequences of length $L=200$ on a binary branching tree with 11 generations and $\mu$ random mutations on each branch, providing $M=2^{11}=2048$ sequences. Finally, data sets with phylogeny and selection are generated along a binary branching tree with $\mu$ accepted mutations per branch (with acceptance criterion in Eq 2 using the same $\kappa$ and $\tau^{*}$ as in the no-phylogeny case and as in Fig. 2) and 11 generations again. Insets show a zoom over large eigenvalues. A logarithmic y-scale is used in the center panel for readability.
@@
We consider the eigenvectors associated to the main outliers identified in Fig 3, i.e. the largest eigenvalue for ICOD, and the smallest one for covariance. Furthermore, as we showed that signal about selection also exists at the other end of the spectrum, we investigate it too. In addition to ICOD and covariance, we consider conservation as a baseline, since it is known to be efficient at identifying sites under selection in natural protein sequence data. How well do these different eigenvectors, or conservation, recover $\vec{D}$ when varying the amount of phylogeny? Fig 4 shows the recovery versus the number of mutations per branch $\mu$ for ICOD, covariance, and conservation. We observe that the eigenvector associated to the largest eigenvalue $\Lambda_{max}$ of ICOD, and the one associated to the smallest eigenvalue $\lambda_{min}$ of covariance (blue), contain more information than those at the opposite end of the respective spectra (purple). This is expected from our theoretical analysis, as the eigenvector with smallest variance should be close to $\vec{D}$ (see above and [31]). Moreover, we observe that ICOD is more robust than covariance to phylogeny, and performs much better than covariance in strong phylogenetic regimes (small $\mu$). Both methods have recoveries close to the maximum possible value of 1 for intermediate to weak phylogeny (larger $\mu$), meaning that they recover almost perfectly $\vec{D}$. Furthermore, ICOD outperforms conservation, except for very small $\mu$ (very strong phylogeny). Conservation performs better than covariance with strong phylogeny, while the opposite holds with weaker phylogeny. We note that the eigenvector associated to the smallest ICOD eigenvalue $\Lambda_{min}$ also contains information about $\vec{D}$, as it outperforms the null model corresponding to recovery by a random vector (see Methods and [31]). The corresponding eigenvector of the covariance is less informative.
@@
Figure 4: Impact of phylogeny on mutational effect recovery. The recovery of the mutational effect vector $\vec{D}$ (see Methods) for specific eigenvectors is shown as a function of the number $\mu$ of mutations per branch, using ICOD and covariance. Recovery by conservation is also shown. For ICOD (resp. covariance $C$), eigenvectors associated to the largest eigenvalue $\Lambda_{max}$ (resp. $\lambda_{max}$) and smallest eigenvalue $\Lambda_{min}$ (resp. $\lambda_{min}$) are considered. Data was generated as in Fig 3, with various values of $\mu$. All results are averaged over 100 realisations of data generation. The null model corresponds to recovery from a random vector (see Methods, Eq 13).
How do these results obtained by ICOD compare to SCA? Recoveries of $\vec{D}$ from SCA are shown in Fig S3, and compared to those of ICOD and conservation. We observe that ICOD is more robust to phylogeny than SCA and generally performs better. Interestingly, we see that conservation yields better recoveries than SCA in strong phylogenetic regimes, while the opposite holds for intermediate to weak phylogeny, where SCA reaches higher recovery scores. Recall that the SCA matrix is constructed using both conservation and covariance.
@@
Impact of selection parameters on mutational effect recovery
Impact of the favored trait value.
To what extent is recovery impacted by the favored trait value $\tau^{*}$, at various levels of phylogeny? To investigate this, we generate data using our minimal model for various values of $\tau^{*}$. Results are presented in Fig 5 for ICOD, covariance, SCA and conservation for three data sets with various levels of phylogeny. First, we observe that ICOD and covariance perform well at small $\tau^{*}$, and then recovery decays and gets poor (below null model) around $\tau^{*}=\sum_{i}D_{i}(\approx 196)$ before it gets somewhat better for ICOD at large $\tau^{*}$. SCA performs rather well at small and intermediate $\tau^{*}$ (except $\tau^{*}=0$) but its performance deteriorates as $\tau^{*}$ increases. Finally, conservation performs well overall and reaches a maximum around $\tau^{*}=\sum_{i}D_{i}(\approx 196)$. Moreover, we observe that ICOD and conservation are more robust to phylogeny than covariance and SCA.
Figure 5: Impact of the favored trait value on mutational effect recovery. The mutational effect recovery is shown as a function of the favored trait value $\tau^{*}$ for ICOD and covariance (left panel), and for SCA and conservation (right panel). As in Fig 3, we consider different levels of phylogeny by considering different values of $\mu$ (shown as different colors). For ICOD (resp. SCA), eigenvectors associated to the largest eigenvalue $\Lambda_{max}$ (resp. $\lambda_{max}$) are considered, while for covariance $C$, the eigenvector associated to the smallest eigenvalue $\lambda_{min}$ is considered. The null model corresponds to recovery from a random vector (see Methods, Eq 13). The data is generated exactly as in Fig 3, except that $\tau^{*}$ is varied. All results are averaged over 100 realisations. Results for $\tau^{*}\geq 0$ are shown, and those for $\tau^{*}<0$ can be obtained by symmetry, see Eq 1.
In Fig S6, we use symmetrized AUC (area under the receiver operating curve, see methods) to assess performance of sector site identification on the same data with the same methods. While ICOD and covariance display similar behaviours with symmetrized AUC as with recovery, SCA and conservation reach better performance with symmetrized AUC, especially for large $\tau^{*}$ or with little phylogeny. Indeed, conservation identifies sites associated to the sector as highly conserved, but does not recover mutational effect values, and SCA may inherit this property from conservation. However, SCA is quite hindered by phylogeny at small $\tau^{*}$, which limits the interest of the contribution of conservation.
Since we showed that the other end of the ICOD spectrum, corresponding to small eigenvalues, also contains information on the sector, we also compute recovery and symmetrized AUC from that end of the spectrum for ICOD, covariance and SCA, see Fig S5 and S7. Those results confirm that the smallest ICOD eigenvalue $\Lambda_{min}$ contains some information on the sector, especially at large values of $\tau^{*}$. Meanwhile, results for covariance and SCA are poor for that end of the spectrum.
Fig 5, S6, S5 and S7 feature particular behaviours around $\tau^{*}=190$. We remark that this value of $\tau^{*}$ is close to $\sum_{i}D_{i}\approx 196$, which is the value taken by the trait $\tau$ when all sites have state $+1$. Given that sector sites have much stronger mutational effects than others, reaching this value of the trait essentially requires all sector sites have state $+1$, which makes them strongly conserved. Accordingly, the performance of conservation in Fig 5 peaks around $\tau^{*}=190$. Indeed, for larger $\tau^{*}$, even non-sector sites have to be conserved, including those with negative mutational effect, impairing recovery but not symmetrized AUC (as sector sites remain the most conserved ones), see Fig S6.
The ICOD spectrum and the corresponding eigenvectors are analyzed in more detail in Fig S9 for different values of $\tau^{*}$. In the spectrum, the number of positive outliers increases when $\tau^{*}$ is increased, and these eigenvalues become very large, while the number of negative outliers reduces to one. For large $\tau^{*}$, sector sites are extremely conserved, and each of them gives rise to a direction of very low variance [31], which results in very large eigenvalues of the ICOD matrix due to the matrix inversion step. In this regime, we do not observe such a split of sector signal at the opposite end of the spectrum (most negative eigenvalue), which contains information on the sector as well (see above). Accordingly, the eigenvector associated to the largest eigenvalue $\Lambda_{max}$ of the ICOD matrix recovers the mutational effect vector $\vec{D}$ very well for moderate $\tau^{*}$, but possesses large components only on a few of the sector sites for larger $\tau^{*}$. Meanwhile, the eigenvector associated to $\Lambda_{min}$ has large components on all sector sites for larger $\tau^{*}$, which is not the case for smaller $\tau^{*}$.
Impact of selection strength on mutational effect recovery.
How do selection strength and phylogeny combine and impact the performance of mutational effect recovery? To address this, we generate data using our minimal model for various values of selection strength $\kappa$. Results are shown in Fig 6 for ICOD and covariance, with two values of $\tau^{*}$ and various levels of phylogeny. For $\tau^{*}=90$, mutational effect recovery by ICOD and covariance is quite robust to selection strength $\kappa$, apart from the very weak selection case where recovery is low because sequences are noisy. Furthermore, covariance is more negatively impacted by phylogeny than ICOD across values of $\kappa$. For $\tau^{*}=140$, we observe that recovery tends to decrease when $\kappa$ increases (ignoring the very weak selection regime). However, this only happens with substantial phylogeny ($\mu=5$) for ICOD, while this effect is always present for covariance, but becomes stronger when phylogeny is stronger. In this case, phylogeny makes recovery by covariance poor, presumably because the combination of strong phylogeny and large $\tau^{*}$ yields strong conservation. Consistently, Fig S8 shows that conservation yields good recovery for $\tau^{*}=140$, which does not deteriorate too much due to phylogeny. The latter figure also shows that recovery by SCA is impaired by strong selection and phylogeny. Overall, ICOD and conservation are the most robust methods to stronger selection and phylogeny.
Figure 6: Impact of selection strength on mutational effect recovery. Recovery by covariance and ICOD is shown as a function of the rescaled selection strength $\tilde{\kappa}=\kappa\sum_{i}D_{i}^{2}$ for three data sets generated with different levels of phylogeny, and with favored trait value $\tau^{*}=90$ (left) or $\tau^{*}=140$ (right). For ICOD, the eigenvector associated to the largest eigenvalue $\Lambda_{max}$ is considered, while for covariance $C$, the eigenvector corresponding to the smallest eigenvalue $\lambda_{min}$ is considered. The data is generated as in Fig 3, except that $\kappa$ is varied. All results are averaged over 100 realisations.
@@
How does phylogeny impact the components of the key eigenvectors of the ICOD, covariance and SCA matrices? To investigate this question, we generate data using our minimal model and keeping track of all intermediate ancestral sequences in the phylogenetic tree. Next, we assign to each site of the sequence a score G, which corresponds to the earliest generation when the site mutates (and thus becomes different from the initial state in the ancestral sequence). In Fig 7, we examine how the eigenvector components relate to G, for two data sets, generated respectively with small and large amounts of phylogeny. In the weak phylogeny regime, ICOD, C and SCA all yield much larger eigenvector components for sector sites than for other sites. Conversely, with strong phylogeny, the eigenvector components associated with sector and non-sector sites strongly overlap when using covariance and SCA. Remarkably, ICOD is still able to disentangle them. We also note that covariance scores are strongly affected by $G$: sites that mutate late have stronger scores, presumably because they are quite conserved and feature little variance, resulting in a high contribution to the eigenvector associated with the smallest eigenvalue of the covariance matrix. Furthermore, Fig S10 considers the eigenvector associated with the largest eigenvalue of the covariance matrix (other end of the spectrum) under strong phylogeny. There, sites that mutate early in the phylogeny obtain high scores. Thus, the spectrum of the covariance matrix is strongly impacted by phylogeny. Similarly, SCA attributes higher scores to sites that mutate relatively early in the phylogeny and might be less conserved, leading to a higher variance and high contributions to the eigenvector associated with the largest eigenvalue of the SCA matrix.

### v2

@@ Abstract
Statistical analysis of multiple sequence alignments of homologous proteins has revealed groups of coevolving amino acids called sectors. These groups of amino-acid sites feature collective correlations in their amino-acid usage, and they are associated to functional properties. Modeling showed that nonlinear selection on an additive functional trait of a protein is generically expected to give rise to a functional sector. These modeling results motivated a principled method, called ICOD, which is designed to identify functional sectors, as well as mutational effects, from sequence data. However, a challenge for all methods aiming to identify sectors from multiple sequence alignments is that correlations in amino-acid usage can also arise from the mere fact that homologous sequences share common ancestry, i.e. from phylogeny. Here, we generate controlled synthetic data from a minimal model comprising both phylogeny and functional sectors. We use this data to dissect the impact of phylogeny on sector identification and on mutational effect inference by different methods. We find that ICOD is most robust to phylogeny, but that conservation is also quite robust. Next, we consider natural multiple sequence alignments of protein families for which deep mutational scan experimental data is available. We show that in this natural data, conservation and ICOD best identify sites with strong functional roles, in agreement with our results on synthetic data. Importantly, these two methods have different premises, since they respectively focus on conservation and on correlations. Thus, their joint use can reveal complementary information.
@@ Introduction
Statistical inference from genome and protein sequence data is currently making great progress. This is due both to the important growth of available genome sequences, and to the development of inference methods ranging from interpretable models inspired by statistical physics and information theory to large deep learning models. One important type of data sets consists in multiple sequence alignments (MSAs) of homologous proteins. These proteins, which are said to be the members of a protein family, share a common ancestry and common functional properties. Because of this, MSA columns (i.e. residue sites) feature correlations in amino-acid usage. Pairwise correlations due to contacts between amino acids in the three-dimensional structure of proteins have been thoroughly studied [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23]. Statistical analyses of MSAs have also revealed the existence of groups of collectively correlated amino acids, termed sectors, which are associated to a functional role and are often spatially close in the protein structure [24, 25, 26, 27, 28, 29]. Preserving the pairwise correlations of natural sequences in a sector-based analysis enabled the successful design of new functional sequences in a pioneering work [26]. Sectors can be identified using Statistical Coupling Analysis (SCA), which focuses on the large-eigenvalue modes of a covariance matrix weighted by conservation [27, 30]. It was recently shown in a minimal model that sectors of collectively correlated amino acids emerge from nonlinear selection on an additive trait (i.e. functional property) of the protein [31]. This case should be quite generic, since empirical data can often be modeled as nonlinear functions of a linear trait [32], and since many protein traits involve additive contributions from amino acids [33, 34, 32]. These modeling results motivated a principled method, called ICOD (for Inverse Covariance Off-Diagonal), which is designed to identify functional sectors, and more generally mutational effects, from sequence data [31]. Here, the focus is on the large-eigenvalue modes of the ICOD matrix, which essentially correspond to the small-eigenvalue modes of the covariance matrix, making ICOD a priori quite distinct from SCA. Another important difference is that ICOD was designed to focus on correlations and eliminate conservation signal as well as possible. Accordingly, a strength of ICOD is its robustness to conservation [31].
@@ Results
To incorporate phylogeny, we take an equilibrium sequence and we evolve this ancestral sequence along a perfect binary tree, see Fig 1B. Along the tree, random mutations changing the state of an amino acids are proposed, and they are accepted or rejected according to the same Metropolis criterion as the one used to generate equilibrium sequences. Thus, selection on the trait is maintained through the phylogeny. A fixed number $\mu$ of accepted mutations are performed on each branch. As a result, two closest (resp. farthest) sequences on the leaves of a tree differ at most by $2\mu$ mutations (resp. $2\mu n$ mutations, where $n$ is the number of generations, i.e. of branching events, in the tree). Note that these numbers are reduced if multiple substitutions occur at the same site. Generated sequences at the leaves of the tree therefore contain correlations from phylogeny and from selection. Their importance are controlled respectively by the parameters $\mu$ and $\kappa$: a smaller $\mu$ means that sequences are more closely related, yielding more phylogenetic correlations, while a larger $\kappa$ means stronger selection, yielding more correlations arising from the sector. Note that this data generation process with phylogeny is close to the one we used previously [47, 40], but that the Hamiltonian we use here (Eq 1) is specific to the sector model.
Figure 1: Data generation process. A: Generation of sequences with selection only. Given a vector of mutational effects $\vec{D}$ on the trait of interest, we sample independent equilibrium sequences under the Hamiltonian $H(\vec{\sigma})$ in Eq 1. For this, we start from random sequences and we use a Metropolis Monte Carlo algorithm where proposed mutations (changes of state at a randomly chosen site) are accepted with probability $p$, according to the Metropolis criterion associated to $H(\vec{\sigma})$. The obtained sequences feature pairwise correlations and conservation arising from selection on the trait (via the Hamiltonian $H(\vec{\sigma})$ in Eq 1). B: To incorporate phylogenetic correlations, we start from one equilibrium sequence, which becomes the ancestor. We evolve it on a perfect binary tree over a fixed number $n$ of “generations” (i.e., tree levels, corresponding to branching events), here $n=2$ generations. A fixed number of mutations $\mu$ (red, here $\mu=2$) are accepted with probability $p$ on each branch of the tree. The earliest generation at which a site mutates with respect to its ancestral state is denoted by $G$, see examples highlighted in green.
@@
Before studying the effect of phylogeny, let us analyze the ICOD matrix and its spectrum without phylogeny, which will serve as baseline. To first order in $\kappa$ (small selection regime), the off-diagonal elements of the ICOD matrix can be approximated as $\tilde{C}_{ij}^{-1}\approx\kappa D_{i}D_{j}$ for all $i,j$ between 1 and $L$ with $i\neq j$ [31]. Here, we show using two different methods that this formula still holds beyond first order in $\kappa$: one method extents this formula to second order, and the second one to fourth order (see Supplementary Appendix, section S1). Within this robust approximation, if diagonal elements were equal to $\kappa D_{i}^{2}$ instead of 0, the spectrum of the ICOD matrix would comprise a single nonzero eigenvalue $\kappa\sum_{i=1}^{L}D_{i}^{2}$ associated to the mutational effect vector $\vec{D}$. Hence, the eigenvector associated to the largest eigenvalue of the ICOD matrix allows to recover $\vec{D}$, as observed in [31].
@@
Note that in the example shown in Fig 2, important mutational effects corresponding to the sector are all positive. We checked that our results are robust to the case where the sector comprises both sites with large positive and and large negative mutational effects: as shown in Fig S2, the spectrum is then very similar to that in Fig 2. Furthermore, the top eigenvector of the ICOD matrix successfully recovers $\vec{D}$.
Figure 2: Spectrum of the ICOD matrix and of its block diagonal approximation. Left (resp. middle) panel: spectrum of the ICOD matrix $\tilde{C}^{-1}$ computed on 2000 (resp. 14,000) sequences generated independently at equilibrium, and of its block diagonal approximation. The spectrum of the inverse covariance matrix $C^{-1}$ is also shown as a reference. Right panel: spectrum of the analytical approximation of the ICOD matrix, and of its block diagonal approximation. Sequences of length $L=200$ were sampled independently at equilibrium using the Hamiltonian in Eq 1 with $\kappa=10/\left(\sum_{i}{D_{i}^{2}}\right)$ and $\tau^{*}=90$. The vector of mutational effect $\vec{D}$ comprises sector sites (the 20 first sites) with components sampled from a Gaussian distribution with mean 5 and variance 0.25, and non-sector sites (the remaining 180 sites) with components sampled from a Gaussian distribution with mean 0.5 and variance 0.25. The analytical approximation $\tilde{C}^{-1}_{ij}\approx(1-\delta_{ij})\kappa D_{i}D_{j}$ (see Supplementary Appendix, section S1) was computed from the values of $\kappa$ and $\vec{D}$ used for data generation.
@@
Figure 3: Impact of phylogeny and selection on ICOD, covariance and SCA spectra. Eigenvalues of the ICOD, covariance and SCA matrices, sorted from largest to smallest, are shown for sequences generated with only phylogeny (light shades) and both phylogeny and selection (dark shades). We consider different levels of phylogeny by considering different values of $\mu$ (shown as different colors). ‘No phylogeny’ corresponds to sequences generated independently at equilibrium, and thus containing only correlations due to selection. This data set comprises $M=2048$ sequences of length $L=200$ generated exactly as in Fig 2, i.e. using the Hamiltonian in Eq 1 with $\kappa=10/\left(\sum_{i}{D_{i}^{2}}\right)$, $\tau^{*}=90$, and the same vector of mutational effect $\vec{D}$ as in Fig 2. Data sets without selection are generated by evolving random sequences of length $L=200$ on a perfect binary branching with 11 generations and $\mu$ random mutations on each branch, providing $M=2^{11}=2048$ sequences. Finally, data sets with phylogeny and selection are generated along a perfect binary tree with $\mu$ accepted mutations per branch (with acceptance criterion in Eq 2 using the same $\kappa$ and $\tau^{*}$ as in the no-phylogeny case and as in Fig 2) and 11 generations again. The three values of $\mu$ shown here were chosen to illustrate different levels of phylogenetic impact. Insets show a zoom over large eigenvalues. A logarithmic y-scale is used in the center panel for readability.
@@
We consider the eigenvectors associated to the main outliers identified in Fig 3, i.e. the largest eigenvalue for ICOD and SCA, and the smallest one for covariance. We also consider conservation as a baseline, since it is known to be efficient at identifying sites under selection in natural protein sequence data (see e.g. [48]). How well do these different eigenvectors, or conservation, recover $\vec{D}$ when varying the amount of phylogeny? Fig 4 shows the recovery versus the number of mutations per branch $\mu$ for ICOD, SCA, covariance, and conservation. We observe that ICOD is more robust than covariance and SCA to phylogeny, and performs much better than covariance in strong phylogenetic regimes (small $\mu$). ICOD and covariance have recoveries close to the maximum possible value of 1 for intermediate to weak phylogeny (larger $\mu$), meaning that they recover almost perfectly $\vec{D}$. Still for large $\mu$, SCA and conservation reach similar performance, which is less good than that of ICOD and covariance. Furthermore, ICOD outperforms conservation, except for very small $\mu$ (very strong phylogeny). Conservation performs better than covariance and SCA with strong phylogeny, while the opposite holds with weaker phylogeny. In Fig 4 and throughout, we employ a small pseudocount for ICOD but not for other approaches (see Methods). This is motivated by the need to invert the covariance matrix in ICOD [31]. For covariance, the difference in the recovery between results with and without pseudocount (taking the same pseudocount value as for ICOD) is lower than 1.2% for all values of $\mu$. For SCA, it is lower than 0.3%, and for conservation it is lower than $3\times 10^{-4}$%. Thus, the pseudocount does not affect the conclusions.
As we showed above that signal about selection also exists at the other end of the spectrum, we investigate it too in Fig S3. The resulting performance is substantially worse than in Fig 4. For ICOD and covariance, this is expected from our theoretical analysis, since it is the eigenvector with smallest variance that should be close to $\vec{D}$ (see above and [31]). Nevertheless, we note that the eigenvector associated to the smallest ICOD eigenvalue $\Lambda_{min}$ also contains information about $\vec{D}$. Indeed, Fig S3 shows that it outperforms the null model corresponding to recovery by a random vector (see Methods and [31]).
@@
Figure 4: Impact of phylogeny on mutational effect recovery. The recovery of the mutational effect vector $\vec{D}$ (see Methods) for specific eigenvectors is shown as a function of the number $\mu$ of mutations per branch, using ICOD, covariance, SCA and conservation. For ICOD (resp. SCA), eigenvectors associated to the largest eigenvalue $\Lambda_{max}$ (resp. $\lambda_{max}$) are considered. For covariance $C$, eigenvectors associated to the smallest eigenvalue $\lambda_{min}$ are considered. Datasets of $M=2048$ sequences of length $L=200$ were generated along a perfect binary tree with 11 generations, using various numbers $\mu$ of accepted mutations per branch. As in Fig 3, we employed the mutation acceptance criterion in Eq 2 with $\kappa=10/\left(\sum_{i}{D_{i}^{2}}\right)$ and $\tau^{*}=90$. We used the same vector of mutational effect $\vec{D}$ as in Fig 2 and Fig 3. All results are averaged over 100 realisations of data generation. The null model corresponds to recovery from a random vector (see Methods, Eq 13).
@@
So far, we discussed the impact of phylogeny on mutational effect recovery for fixed values of the favored trait value $\tau^{*}$ and of the selection strength $\kappa$. These two parameters, which characterize selection, impact recovery too [31]. We study their interplay with phylogeny in section S2 of the Supplementary Appendix.
Generalization to other phylogenetic trees.
How does the type of phylogenetic tree affect the performance of sector inference? So far, and in particular in Fig 4, synthetic data was generated using a perfect binary tree for simplicity (see Methods). However, natural phylogenies are more complex. To assess the impact of more general phylogenetic trees on our results, we generate synthetic data using Beta-coalescent trees, following Ref. [49], instead of perfect binary trees. These trees range from a Bolthausen-Snitzman coalescent (with selection) for a Beta-coalescent tree with $\alpha=1$ to a Kingman coalescent (neutral case) for a Beta-coalescent tree with $\alpha=2$, see examples in Fig S5. In Fig S6, we show the impact of branch length on mutational effect recovery, as in Fig 4, but for the trees shown in Fig S5. Throughout Fig S6, we find similar results regarding the relative performance of ICOD, conservation and covariance as in Fig 4. Furthermore, our results regarding the performance of SCA compared to other methods are also similar to those in Fig 4 for $\alpha=1$ and $\alpha=1.2$. For $\alpha=1.4$ and $\alpha=1.6$, we observe that, in one tree realization out of two, SCA has lower performance with strong phylogeny than for other trees, and features a slower convergence to its no-phylogeny asymptote. Conversely, for $\alpha=1.8$ and $\alpha=2$, we find that SCA performs better than in other cases, and reaches performances similar or slightly better than those of other methods. These results suggest that SCA is more sensitive than other methods to the details of a phylogenetic tree, and may be particularly useful for neutral or quasi-neutral cases. Apart from these points, our main results are robust to considering diverse phylogenetic trees.
@@
How does phylogeny impact the components of the key eigenvectors of the ICOD, covariance and SCA matrices? To investigate this question, we generate data using our minimal model and keeping track of all intermediate ancestral sequences in the phylogenetic tree. Next, we assign to each site of the sequence a score $G$, which corresponds to the earliest generation, i.e., tree level, at which the site mutates, and thus becomes different from the initial state in the ancestral sequence, see Fig 1B. In Fig 5, we examine how the eigenvector components relate to $G$, for two data sets, generated with two different values of the number $\mu$ of mutations per branch of the tree. These values yield respectively small and large effects of phylogeny on the mutational effect recovery in Fig 4, and match two of the three values shown in Fig 3. The mean pairwise Hamming distances in these synthetic datasets are 0.47 for $\mu=50$ and 0.30 for $\mu=5$. They fall in the range observed in natural data, see Tables S1 and S2. Fig 5 shows that in the weak phylogeny regime, ICOD, covariance and SCA all yield much larger eigenvector components for sector sites than for other sites. Conversely, with strong phylogeny, the eigenvector components associated with sector and non-sector sites strongly overlap when using covariance and SCA. Remarkably, ICOD is still able to disentangle them. We also note that covariance scores are strongly affected by $G$: sites that mutate late have stronger scores, presumably because they are quite conserved and thus feature little variance, resulting in a high contribution to the eigenvector associated with the smallest eigenvalue of the covariance matrix. Recall that while the eigenvector associated to the largest ICOD eigenvalue also focuses on small variances, the diagonal of the ICOD matrix is set to 0 in order to remove conservation effects. Consistently, the contributions of sites to the ICOD eigenvector only increase moderately with $G$. Furthermore, Fig S7 considers the eigenvector associated with the largest eigenvalue of the covariance matrix (other end of the spectrum) under strong phylogeny. There, sites that mutate early in the phylogeny obtain high scores. Thus, the spectrum of the covariance matrix is strongly impacted by phylogeny. Similarly, SCA attributes higher scores to sites that mutate relatively early in the phylogeny and might be less conserved, leading to a higher variance and high contributions to the eigenvector associated with the largest eigenvalue of the SCA matrix.